OpenStudy (anonymous):

Solve: 4(x + 1) = −4

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

4x+4=-4 =>4x=-4-4 =>4x=-8 x=-2

hero (hero):

Divide both sides by 4 to get x + 1 = -1 x = -1 - 1 x = -2

OpenStudy (anonymous):

ty

hero (hero):

Dividing keeps the computation in a most simplest form. Helps to avoid having to deal with bigger numbers.